### v2

# Graham’s rearrangement for a class of semidirect products

Abstract.
A famous conjecture of Graham asserts that every set $A\subseteq\mathbb{Z}_{p}\setminus\{0\}$ can be ordered so that all partial sums are distinct. Bedert and Kravitz proved in [4] that this statement holds whenever $|A|\leq e^{c(\log p)^{1/4}}$.
In this paper, we will use a similar procedure to obtain an upper bound of the same type in the case of semidirect products $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$ and where $H$ is abelian and each subset of $H$ can be ordered such that all of its partial products are distinct.
Key words and phrases:
Sequenceability, Rectification
1991 Mathematics Subject Classification:
11B75
1. Introduction
Let $A$ be a finite subset of group $(G,\cdot)$. We say that an ordering $a_{1},\ldots,a_{|A|}$ of $A$ is valid if the partial products (or partial sums in additive notation) $p_{1}=a_{1},p_{2}=a_{1}\cdot a_{2},\ldots,p_{|A|}=a_{1}\cdot a_{2}\cdot\cdots\cdot a_{|A|}$ are all distinct. Moreover, this ordering is a sequencing if it is valid and $p_{i}\not=id$ for any $i\in[1,|A|-1]$. In this case, we say that $A$ is sequenceable. We also say that a group $G$ is strongly sequenceable if all the subsets of $G\setminus\{id\}$ are sequenceable. In the literature, there are several conjectures about valid orderings and sequenceability. We refer to [8, 12, 14] for an overview of the topic, [1, 2, 3, 7] for lists of related conjectures, and [5] for a treatment using rainbow paths. Here, we explicitly recall only the conjecture of Graham.
Conjecture 1.1 ([10] and [9]).
Let $p$ be a prime. Then every subset $A\subseteq\mathbb{Z}_{p}\setminus\{0\}$ has a valid ordering.
Until recently, the main results on this conjecture were for small values of $|A|$; in particular, the conjecture was proved for sets $A$ of size at most $12$ [7].
The first result involving arbitrarily large sets $A$ was presented last year by Kravitz [11]. He used a simple rectification argument to show that Graham’s conjecture holds for all sets $A$ of size at most $\log p/\log\log p$. A similar argument was also proposed (but not published) by Sawin [15] in a 2015 MathOverflow post.
Then, in [4], Bedert and Kravitz improved this upper bound to the following:
Theorem 1.2 ([4]).
Let $c>0$ be a constant and $p$ be a large enough prime. Then every subset $A\subseteq\mathbb{Z}_{p}\setminus\{0\}$ is sequenceable provided that
$|A|\leq e^{c(\log p)^{1/4}}$
Graham’s conjecture over non-abelian groups was first studied by Ollis in [13] over dihedral groups. More recently, the authors in [5] presented a nice connection with rainbow paths and produced the asymptotic result that any subset $A$ of a finite group $G$ admits an ordering in which at least (1o(1))|A|(1-o(1))|A| many partial products are distinct. Inspired by the result of Kravitz ([11]) on cyclic groups, Costa and Della Fiore proved in [6] that any subset $A$ of $Dih_{p}\setminus\{id\}$, where $p$ is prime and $Dih_{p}$ is the dihedral group of size $2p$, and in general any subset $A$ of a class of semidirect products, is sequenceable provided that $p>|A|!/2$. Note that this is slightly better than $|A|\leq\log p/\log\log p$.
Section 3 of this paper is dedicated to proving, with a similar procedure, the sequenceability of subsets of semidirect products $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$ and where $H$ is abelian and strongly sequenceable. We note that for $H=\mathbb{Z}_{2}$, we are considering the dihedral group $Dih_{p}$. In this setting, we improve the results of [6] on dihedral groups, obtaining essentially the same bound as in the case of $\mathbb{Z}_{p}$ (with a different best possible constant $c$). That is,
Theorem 1.3.
Let $H$ be a finite abelian and strongly sequenceable group. There exists a $c>0$ such that every subset $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$, where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$, of size
$|A|\leq e^{c(\log p)^{1/4}},$
has a sequencing.
The main difficulty in adapting the approach of [4] to get this theorem is that the semidirect products are not necessary abelian and hence the product of the elements of a set is not fixed. For this reason, we have to impose some restrictions on the positions of certain elements.
2. Proof Strategy
We offer a brief summary of the main steps of our proof of Theorem 1.3. Note that we represent elements in $\mathbb{Z}_{p}\rtimes_{\varphi}H$ as $(x_{i},a_{i})$ where $x_{i}\in\mathbb{Z}_{p}$ and $a_{i}\in H$.
We begin Section 3 by describing how a small set $B\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ can be “rectified” such that all of its $\mathbb{Z}_{p}$ components are contained within a small interval. Then, for a nonempty subset $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$, we prove the existence of an automorphism $\phi\in Aut(\mathbb{Z}_{p}\rtimes_{\varphi}H)$ such that $\phi(A)=E\cup\left(\cup_{j=1}^{s}D_{j}\right)$, where each $D_{j}$ is a dissociated set with size a multiple of eight and where the $\mathbb{Z}_{p}$ components of $E$ are in a small interval by our rectification argument. Moreover, we can impose that each dissociated set $D_{j}$ has elements $d=(x,a)\in D_{j}$ such that either $\varphi(a)=-id$ for all $a$ (in which case we say that $j\in L_{1}$) or $\varphi(a)=id$ for all $a$ ($j\in L_{0}$). Here we split each $D_{j}$ with $j\in L_{1}$ into an odd-position subset $D^{o}_{j}$ and an even-position subset $D^{e}_{j}$ of about equal sizes. Then the product $\tau_{j}$ of the elements of $D_{j}$, with elements of $D^{o}_{j}$ in odd position and elements of $D^{e}_{j}$ in even position, does not depend on the orderings within $D_{j}^{o}$ and $D_{j}^{e}$. This means that, even if $\mathbb{Z}_{p}\rtimes_{\varphi}H$ is non-abelian, we have a constant product $\delta$ of the elements of the dissociated sets without fixing the ordering of all the elements. We then relabel $\phi(A)$ by AA for ease of notation.
In Section 4, we order the elements of $E$. Here we split $E$ into sets $P$ (with $\mathbb{Z}_{p}$ components positive [^1] Here we say the positive elements are the ones in the interval $1,\dots,(p-1)/2]$ while the negatives in [(p1)/2,1][-(p-1)/2,-1]. and $H$ components such that $\varphi(a)=id$), NN (with $\mathbb{Z}_{p}$ components negative and $H$ components such that $\varphi(a)=id$), $Z$ (with $\mathbb{Z}_{p}$ components always zero and $H$ components such that $\varphi(a)=id$), and $S$ (with $H$ components such that $\varphi(a)=-id$). We further split $S$ into subsets $S_{e}$ and $S_{o}$ of almost equal size, where the $\mathbb{Z}_{p}$ component of each element in $S_{e}$ is larger than the $\mathbb{Z}_{p}$ components of all of the elements in $S_{o}$. We then prove the existence of an ordering $\mathbf{x}$ of the elements of $P,N,Z,S_{e},$ and $S_{o}$ such that $\delta,\mathbf{x}$ is a sequencing, where $\delta$ is the prespecified product of the elements of the $D_{j}$s. At the same time, we derive an important bound on the size of overlap between subsets of $\mathbb{Z}_{p}\rtimes_{\varphi}H$ and the set $\delta$ times the partial products of $\mathbf{x}$}.
Finally, we describe a two-step method of ordering the $D_{j}$s. First, we randomly split each $D^{o}_{j}$, each $D^{e}_{j}$ (when $j\in L_{1}$), and each remaining $D_{j}$ (when $j\in L_{0}$) into a four-set partition. Here, we use a probabilistic argument to prove that, given a sequencing $\mathbf{x}$ of $E$, we can obtain a sequencing $\tau_{1},\dots\tau_{u},\mathbf{x}$ of $E$ and of the products $\tau_{i}$ of the partitions of the $D_{j}$s. To conclude, we use again a probabilistic argument to prove that if $\mathbf{x}$ satisfies a technical condition on the intersection between suitable subsets and the set $\delta$ times the partial products of $\mathbf{x}$}, then there exists a sequencing of $E$ and each partition $T_{i}$ of the $D_{j}$s. This gives us Theorem 1.3.
3. Semidirect products
In this section, we consider groups that are semidirect products of the form $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ for some group homomorphism $\varphi:H\rightarrow Aut(\mathbb{Z}_{p})$. We recall that the group operation is defined as
$(x_{1},a_{1})\cdot_{G}(x_{2},a_{2})=(x_{1}+\varphi(a_{1})(x_{2}),a_{1}\cdot_{H}a_{2}).$
Here the map
$x\rightarrow\varphi(a_{1})(x)$
is an automorphism of $(\mathbb{Z}_{p},+)$. It is well known that the automorphisms of $(\mathbb{Z}_{p},+)$ are the multiplications and hence that there exists $\varphi_{a_{1}}\in\mathbb{Z}_{p}$ such that
$\varphi(a_{1})(x)=\varphi_{a_{1}}x.$
Thus for this kind of semidirect product, we can rewrite the group operation as
$(x_{1},a_{1})\cdot_{G}(x_{2},a_{2})=(x_{1}+\varphi(a_{1})(x_{2}))\cdot_{G}(a_{1},a_{2})=(x_{1}+(\varphi_{a_{1}}x_{2}),a_{1}\cdot_{H}a_{2}).$
Here we let $Dih_{p}$ denote the dihedral group $\mathbb{Z}_{p}\rtimes_{\varphi}\mathbb{Z}_{2}$, where $\varphi(1)(x)=-x$ and $\varphi(0)(x)=x$. In this case, the group operation is defined as
$(x_{1},a_{1})\cdot_{G}(x_{2},a_{2})=(x_{1}+(\varphi_{a_{1}}x_{2}),a_{1}+a_{2})$
where $\varphi_{0}=1$ and $\varphi_{1}=-1$. If we set $id:\mathbb{Z}_{p}\rightarrow\mathbb{Z}_{p}$ to be the map id(x)=x-id(x)=-x, then dihedral groups are an example of a semidirect product $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi:H\rightarrow Aut(\mathbb{Z}_{p})$ satisfies the following property:
$(\star)\qquad\qquad\varphi(h)\in\{id,-id\}.$
In the following, we consider only semidirect products that satisfy property $(\star)$.
Given a not necessarily abelian group $G$, a subset $D=\{d_{1},\ldots,d_{r}\}\subseteq G$ is dissociated if
$d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}\neq id$
for all permutations $\sigma\in Sym(r)$ and $(\epsilon_{\sigma(1)},\ldots,\epsilon_{\sigma(r)})\in\{-1,0,1\}^{r}\setminus\{(0,\ldots,0)\}$.
The dimension of a subset $B\subseteq G$, written $\dim(B)$, is the size of the largest dissociated set contained in $B$. We can characterize these maximal dissociated subsets as follows:
Lemma 3.1.
Let $B\subseteq G$ be a finite subset of a not necessarily abelian group. If $D=\{d_{1},\ldots,d_{r}\}$ is a maximal dissociated subset of $B$, then
$B\subseteq\operatorname{span}(D)\vcentcolon=\left\{d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}:\sigma\in Sym(r),\epsilon_{\sigma(i)}\in\{-1,0,1\}\right\}.$
Proof.
Due to the maximality of $D$, for every element $b\in B\setminus D$, $\{b\}\cup D$ is not dissociated. This means that there exists an ordering $\sigma$ and coefficients $\epsilon_{\sigma(i)}\in\{-1,0,1\}$ such that
$d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots d_{\sigma(i-1)}^{\epsilon_{\sigma(i-1)}}\cdot b\cdot d_{\sigma(i)}^{\epsilon_{\sigma(i)}}\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}=id.$
It follows that we have
$b=\left(d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots d_{\sigma(i-1)}^{\epsilon{\sigma(i-1)}}\right)^{-1}\left(d_{\sigma(i)}^{\epsilon(i)}\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}\right)^{-1}$
and hence $b\in Span(D)$. ∎
Next, the following lemma relates the size of a set and its dimension.
Lemma 3.2.
Let $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian and $\varphi$ satisfies property $(\star)$. Let $B$ be a finite subset of $G$. Then $|B|\leq 5^{\dim(B)}$.
Proof.
Let $D=\{(x_{1},a_{1}),(x_{2},a_{2}),\dots,(x_{r},a_{r})\}$ be a maximal dissociated subset of $B$ of size $dim(B)$. Due to the previous lemma, this means that any element $(x,a)$ of $B$ belongs to $\operatorname{span}(D)$ and hence
$(x,a)=d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots\cdot d_{\sigma(t)}^{\epsilon_{\sigma(t)}}$
where $d_{i}=(x_{i},a_{i})\in D$, $\sigma\in Sym(r),\epsilon_{\sigma(1)},\dots,\epsilon_{\sigma(t)}\in\{-1,1\},$ and $\epsilon_{\sigma(t+1)}=\cdots=\epsilon_{\sigma(r)}=0.$ In components, this means
$(x,a)$ $=\left(\pm x_{\sigma(1)}\pm\varphi_{a_{\sigma(1)}}x_{\sigma(2)}+\dots\pm\varphi_{a_{\sigma(1)}+\cdots+a_{\sigma(t-1)}}x_{\sigma(t)},\sum_{i=1}^{t}\pm a_{\sigma(i)}\right)$
$=\left(\gamma_{1}x_{1}+\gamma_{2}x_{2}+\dots+\gamma_{r}x_{r},\lambda_{1}a_{1}+\lambda_{2}a_{2}+\dots+\lambda_{r}a_{r}\right)$
where $(\gamma_{i},\lambda_{i})\in\{(0,0),(1,1),(-1,1),(1,-1),(-1,-1)\}$. Since we have five choices for each pair $(\gamma_{i},\lambda_{i})$, it follows that $|B|\leq 5^{\dim(B)}$. ∎
Following [4], we state a lemma that allows us to “rectify” small sets. We define the parameter
(1) $R=R(A):=c_{1}\max\left((\log p)^{1/2},\frac{\log p}{\log|A|}\right),$
where $c_{1}$ is a sufficiently small absolute constant.
Lemma 3.3.
Let $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian and $\varphi$ satisfies property $(\star)$. If $B\subseteq G$ is a nonempty subset of dimension $\dim(B)<R$, then there is some $\phi\in Aut(G)$ such that the image $\pi_{1}(\phi(B))$ (where $\pi_{1}$ is the projection over the $\mathbb{Z}_{p}$ component) is contained in the interval $(-\frac{p}{100|B|},\frac{p}{100|B|})$.
Proof.
Given $\lambda\in\mathbb{Z}_{p}$, we consider here the family of maps $\phi_{\lambda}:G\rightarrow G$ defined as
$\phi_{\lambda}(x,a)=(\lambda x,a).$
Note that, if $\lambda\not=0$, then $\phi_{\lambda}\in Aut(G)$.
Let us now consider a maximal dissociated subset $D=\{(x_{1},a_{1}),\dots,(x_{r},a_{r})\}$ of $B$, where $r=|D|=\dim(B)$. Lemma 3.1 tells us that $B\subseteq\operatorname{span}(D)$.
Consider the set
$\{\lambda(x_{1},\dots,x_{r}):\lambda\in\mathbb{Z}_{p}\}\subseteq(\mathbb{Z}_{p})^{r}.$
Due to the pigeonhole principle, there exists distinct $\lambda_{1},\lambda_{2}\in\mathbb{Z}_{p}$ such that $\|\lambda_{1}x_{i}-\lambda_{2}x_{i}\|\leq p^{1-1/r}$ for all $i\in[1,r]$. Set $\lambda\vcentcolon=\lambda_{1}-\lambda_{2}$, we have that $\phi_{\lambda}\in Aut(G)$ and $\pi_{1}(\phi_{\lambda}(D))\subseteq[-p^{1-1/\dim(B)},p^{1-1/\dim(B)}]$. Since $B\subseteq\operatorname{span}(D)$, we have
$\pi_{1}(\phi(B))\subseteq[-\dim(B)p^{1-1/\dim(B)},\dim(B)p^{1-1/\dim(B)}].$
The thesis follows if we prove that
(2) $\dim(B)p^{1-1/\dim(B)}<p/(100|B|)$
as long as $c_{1}$ is chosen to be sufficiently small.
Set $h=dim(B)$, Equation (2) is equivalent to
$\log({hp^{1-1/h}})=\log(h)+\log{p}-1/h\log{p}<\log{p}-\log(100)-\log{|B|}$
and so to
(3) $h\log(h)+h\log(100)+h\log|B|<\log p.$
We first consider the case where $R=c_{1}(\log p)^{1/2}$. We note that due to Lemma 3.2 this relation is implied by
$h\log{h}+2h\log{10}+h^{2}\log{5}<\log{p}$
which holds since we have assumed that $h<R=c_{1}(\log p)^{1/2}$.
Let know consider the case where $h<R=c_{1}\log p/\log|B|$. Here Equation (3) holds since for $c_{1}$ small enough we have
$1/3\log p\geq\max\left\{h\log h,h\log(100),h\log|B|\right\}.$
Before stating the structure theorem for semidirect products, we prove the following lemma, which allows us to modify a dissociated set by absorbing a given element.
Lemma 3.4.
Let $G$ be a not necessarily abelian group and let $D_{1}\cup D_{2}$ be a partition of a dissociated subset $D$ of $G$. For every element $x\in G\setminus\{id\}$, either $D_{1}\cup\{x\}$ or $D_{2}\cup\{x\}$ is dissociated.
Proof.
Let us set $D_{1}=\{d_{1},\dots,d_{t}\}$ and $D_{2}=\{d_{t+1},\dots,d_{r}\}$. Assume by contradiction that neither $D_{1}\cup\{x\}$ nor $D_{2}\cup\{x\}$ is dissociated. Since $D_{1}$ and $D_{2}$ are dissociated, Lemma 3.1 tells us that
$x\in Span(D_{1})\cap Span(D_{2}).$
Therefore, we can write
$x=d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots\cdot d_{\sigma(t)}^{\epsilon_{\sigma(t)}}=d_{\sigma(t+1)}^{\epsilon_{\sigma(t+1)}}\cdot\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}}$
and hence
$(d_{\sigma(1)}^{\epsilon_{\sigma(1)}}\cdot\cdots\cdot d_{\sigma(t)}^{\epsilon_{\sigma(t)}})(d_{\sigma(t+1)}^{\epsilon_{\sigma(t+1)}}\cdot\cdots\cdot d_{\sigma(r)}^{\epsilon_{\sigma(r)}})^{-1}=id$
which implies the contradiction that $D$ would not be a dissociated set. ∎
Now we show that any subset $A$ of $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$, where $H$ is abelian and $\varphi$ satisfies property $(\star)$, can be split into a family of dissociated sets and a remainder set $E$ of small dimension which can be embedded, as a consequence of the rectification lemma, into $\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$. We assume that $A$ is not contained in $\{(0,h):h\in H\}$ since $H$ is strongly sequenceable and therefore there exists a sequencing of $A$.
Using standard asymptotic notation, we say that $f\asymp g$ if there exists a constant $C>0$ such that $f=Cg$. We also define $f(p)=o(g(p))$ if $\lim_{p\to\infty}f(p)/g(p)=0$.
Theorem 3.5 (Structure Theorem).
Let $p$ be large enough and let $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian, is strongly sequenceable, and $\varphi$ satisfies property $(\star)$. For every nonempty subset $A\subseteq G\setminus\{id\}$, there is some $\phi\in Aut(G)$ such that $\phi(A)$ can be partitioned as
$\phi(A)=E\cup\left(\cup_{j=1}^{s}D_{j}\right),$
where
  1. (i)
    $dim(E)<R$.
Moreover, if $s>0$,
  1. (ii)
    Each $D_{j}$ is a dissociated set of size $|D_{j}|\asymp R$ and $|D_{j}|$ is a multiple of eight.
  2. (iii)
    We can split the interval $1,s]=L_{0}\cup L_{1}$ so that $1$ and $s$ belong to the same set $L_{i}$, and for any $j\in L_{1}$ we have that $D_{j}\subseteq\{(x,a):x\in\mathbb{Z}_{p}\}$ with $\varphi(a)=-id$, and for any $j\in L_{0}$ we have that $D_{j}\subseteq\{(x,b):x\in\mathbb{Z}_{p}\}$ with $\varphi(b)=id$.
  3. (iv)
    Given $j\in L_{1}$, we split $D_{j}$ into $D_{j}^{o}=\{d_{j,1}^{o},\ldots,d_{j,|D_{j}^{o}|}^{o}\}$ and $D_{j}^{e}=\{d_{j,1}^{e},\ldots,d_{j,|D_{j}^{e}|}^{e}\}$ where $|D_{j}^{o}|=|D_{j}^{e}|=|D_{j}|/2$. Then we set
    $\delta\vcentcolon=\left(\prod_{j\in L_{1}}d_{j,1}^{o}\cdot d_{j,1}^{e}\cdot d_{j,2}^{o}\cdot d_{j,2}^{e}\cdots d_{j,|D_{j}^{o}|}^{o}\cdot d_{j,|D_{j}^{e}|}^{e}\right)\left(\prod_{j\in L_{0}}\prod_{d\in D_{j}}d\right)$
    which does not depend [^2] We will use the notation $\prod_{d^{o}\in D^{o},d^{e}\in D^{e}}d^{o}\cdot d^{e}$ without specifying the orderings if the result does not depend on them. on the orderings on $D_{j}^{o}$, $D_{j}^{e}$, $D_{j}$, and on $L_{0}$ and $L_{1}$. Then $\delta=(z_{0},a_{0})$ where $\varphi(a_{0})=id$, $z_{0}\not=0$, and
    $\pi_{1}(E\cup\{\delta\})\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right).$
  4. (v)
    $D_{1}\cup D_{s}\cup\{\delta\}$ is a dissociated set.
Proof.
Following [4], we begin with $E^{\prime}=A$ and assume that $\dim(E^{\prime})>R$ (i.e. $s>0$). Then, as long as $\dim(E^{\prime})\geq R$, we can iteratively remove a dissociated subset $D^{\prime\prime\prime}$ of size $R/2$ and hence one, say $D^{\prime\prime}$, whose size is in the interval [R/431,R/2][R/4-31,R/2] and congruent to one modulo thirty-two, and such that $\pi_{2}(D^{\prime\prime})$ (i.e. the projection over $H$) is constantly in $H_{0}$ (i.e. the set $h\in H:\varphi(h)=id$) or constantly in $H_{1}$ (i.e. the set $h\in H:\varphi(h)=-id$). Also, if we divide $D^{\prime\prime}$ into two sets ${D^{o}}^{\prime\prime}=\{{d_{1}^{o}}^{\prime\prime},{d_{2}^{o}}^{\prime\prime}\dots,{d_{h+1}^{o}}^{\prime\prime}\}$ and ${D^{e}}^{\prime\prime}=\{{d_{1}^{e}}^{\prime\prime},{d_{2}^{e}}^{\prime\prime}\dots,{d_{h}^{e}}^{\prime\prime}\}$, then
$\delta^{\prime\prime}\vcentcolon={d_{1}^{o}}^{\prime\prime}\cdot{d_{1}^{e}}^{\prime\prime}\cdots{d_{h}^{o}}^{\prime\prime}\cdot{d_{h}^{e}}^{\prime\prime}\cdot{d_{h+1}^{o}}^{\prime\prime}$
does not depend on the orderings on $D^{o}$$^{\prime\prime}$ and $D^{e}$$^{\prime\prime}$. Note that up to deleting one element from $D^{o}$$^{\prime\prime}$, we can always assume that $\delta^{\prime\prime}$ is different from a given element $\bar{x}\in G$. This means that as long as $\dim(E^{\prime})\geq R$, we can iteratively remove a dissociated subset $D^{\prime}$, whose size is in the interval [R/432,R/2][R/4-32,R/2] and is a multiple of thirty-two, and such that $\pi_{2}(D^{\prime})$ is constantly in $H_{0}$ or constantly in $H_{1}$ and the product of its elements (in a given order) is different from a chosen element $\bar{x}\in G$.
Once we reach a residual set $E^{\prime}$ of dimension smaller than $R$, we obtain a decomposition
$A=E^{\prime}\cup(\cup_{j\in L_{1}^{\prime}}D_{j}^{\prime})\cup(\cup_{j\in L_{0}^{\prime}}^{s^{\prime}}D_{j}^{\prime}),$
where each $D_{j}^{\prime}$ is a dissociated set whose size is in the interval $R/4-32\leq|D_{j}^{\prime}|\leq R/2$ and is a multiple of thirty-two, for any $j\in L_{1}^{\prime}$ we have that $D^{\prime}_{j}\subseteq\{(x,a):x\in\mathbb{Z}_{p},a\in H_{1}\}$, and for any $j\in L_{0}^{\prime}$ we have that $D^{\prime}_{j}\subseteq\{(x,a):x\in\mathbb{Z}_{p},a\in H_{0}\}$.
Given $j\in L_{1}^{\prime}$, we split $D_{j}^{\prime}$ into $D_{j}^{o}$$^{\prime}$ and $D_{j}^{e}$$^{\prime}$ so that $|{D_{j}^{o}}^{\prime}|=|{D_{j}^{e}}^{\prime}|$. Then we set
$\delta^{\prime}=\left(\prod_{j\in L_{1}^{\prime}}\prod_{{d^{o}}^{\prime}\in{D_{j}^{o}}^{\prime},{d^{e}}^{\prime}\in{D_{j}^{e}}^{\prime}}{d^{o}}^{\prime}\cdot{d^{e}}^{\prime}\right)\left(\prod_{j\in L_{0}^{\prime}}\prod_{d^{\prime}\in D^{\prime}_{j}}d^{\prime}\right),$
which does not depend on the orderings on $D_{j}^{o}$$^{\prime}$ and $D_{j}^{e}$$^{\prime}$. Also, due to the previous consideration, we can assume that $\delta^{\prime}\not=(0,0)$.
Then Lemma 3.3 (applied to $E^{\prime}\cup\{\delta^{\prime}\}$) provides an automorphism $\phi\in Aut(G)$ such that
$\phi(A)=D_{1}\cup\cdots\cup D_{s^{\prime}}\cup E,$
where $E$ satisfies property $(i)$ and each $D_{j}$ is a dissociated set of size a multiple of thirty-two which satisfies properties $(ii)$ (with the stronger lower bound R/432R/4-32) and $(iii)$.
Setting $D_{j}^{o}=\phi({D_{j}^{o}}^{\prime})$, $D_{j}^{e}=\phi({D_{j}^{e}}^{\prime})$, and $\delta=\phi(\delta^{\prime})$, we have that, since $|D_{j}|$ is even for any $j$, $\delta$ is of the form $(z_{0},a_{0})$ where $z_{0}\not=0$ and $a_{0}\in H_{0}$. Also, due to the choice of $\phi$,
$\pi_{1}(\phi(E^{\prime}\cup\{\delta^{\prime}\}))=\pi_{1}(E\cup\{\delta\})\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$
which implies condition $(iv)$.
Following again [4], we can modify the decomposition to satisfy (v) as follows. We split $D_{1}^{o}$ and $D_{1}^{e}$ into $2$ parts $D_{1}^{o,(1)},D_{1}^{o,(2)}$ and $D_{1}^{e,(1)},D_{1}^{e,(2)}$ each of size $|D_{1}|/4\geq R/16-8$. Lemma 3.4 ensures that either $D_{1}^{o,(1)}\cup D_{1}^{e,(1)}\cup\{\delta\}$ or $D_{1}^{o,(2)}\cup D_{1}^{e,(2)}\cup\{\delta\}$ is dissociated. Setting $D_{1}^{(1)}=D_{1}^{o,(1)}\cup D_{1}^{e,(1)}$, without loss of generality, we can assume that $D_{1}^{(1)}\cup\{\delta\}$ is dissociated.
Then further split $D_{1}^{o,(1)}$ and $D_{1}^{e,(1)}$ into $2$ parts $D_{1}^{o,(3)},D_{1}^{o,(4)}$ and $D_{1}^{e,(3)},D_{1}^{e,(4)}$ each of size $|D_{1}|/8\geq R/32-4$. Replace the sequence of sets $D_{1},\ldots,D_{s^{\prime}}$ by the sequence
$D_{1}^{(3)},D_{1}^{(2)},D_{2},D_{3},\ldots,D_{s^{\prime}},D_{1}^{(4)},$
where $D_{1}^{(2)}=D_{1}^{o,(2)}\cup D_{1}^{e,(2)}$, $D_{1}^{(3)}=D_{1}^{o,(3)}\cup D_{1}^{e,(3)}$, and $D_{1}^{(4)}=D_{1}^{o,(4)}\cup D_{1}^{e,(4)}.$ For notational convenience, we rename this sequence as
$D_{1},\ldots,D_{s}.$
Since this procedure does not change the value of $\delta$, the new decomposition satisfies all the required properties. ∎
Note that, in the Structure Theorem 3.5, if $s>0$ then we always obtain a decomposition where $|E|\geq R/2$.
We conclude this section by stating a structural property of dissociated sets that is important for the probabilistic section of this paper.
Lemma 3.6.
Consider a dissociated set $D=D^{o}\cup D^{e}\subseteq G$ where $|D^{o}|=\lceil|D|/2\rceil$ and $|D^{e}|=\lfloor|D|/2\rfloor$ for $G=\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian, $\varphi$ satisfies property $(\star)$, and the $H$ component of each element of $D$ is in $H_{1}$. Then for any
$D_{i}=\{d_{1},\dots,d_{h}\},\ D_{j}=\{d_{1}^{\prime},\dots,d_{\ell}^{\prime}\}\subseteq D$
where $D_{i}\neq D_{j}$, $d_{k},d_{k}^{\prime}\in D^{o}$ if $k$ is odd, and $d_{k},d_{k}^{\prime}\in D^{e}$ if $k$ is even, we have
$d_{1}d_{2}\cdots d_{h}\neq d_{1}^{\prime}d_{2}^{\prime}\cdots d_{\ell}^{\prime}.$
Proof.
Assume for the sake of contradiction that there exist $D_{i}=\{d_{1},\dots,d_{h}\},\ D_{j}=\{d_{1}^{\prime},\dots,d_{\ell}^{\prime}\}\subseteq D$ where $D_{i}\neq D_{j}$, $d_{k},d_{k}^{\prime}\in D^{o}$ if $k$ is odd, $d_{k},d_{k}^{\prime}\in D^{e}$ if $k$ is even, and $d_{1}d_{2}\cdots d_{h}=d_{1}^{\prime}d_{2}^{\prime}\cdots d_{\ell}^{\prime}.$ Setting $d_{k}=(x_{k},a_{k})$ and $d_{k}^{\prime}=(x_{k}^{\prime},a_{k}^{\prime})$, in components we have that
(4) $(x_{1}-x_{2}+x_{3}-\dots(-1)^{h+1}x_{h},\sum_{i=1}^{h}a_{i})=(x_{1}^{\prime}-x_{2}^{\prime}+x_{3}^{\prime}-\dots(-1)^{\ell+1}x_{\ell}^{\prime},\sum_{i=1}^{\ell}a_{i}^{\prime}).$
This is only possible if $h\equiv\ell\pmod{2}$ since $(-id)^{h}=\varphi(\sum_{i=1}^{h}a_{i})=\varphi(\sum_{i=1}^{\ell}a^{\prime}_{i})=(-id)^{\ell}$ and by moving everything to the left-hand side, we get
(5) $(x_{1}-x_{2}+x_{3}-\dots(-1)^{h+1}x_{h}-(x_{1}^{\prime}-x_{2}^{\prime}+x_{3}^{\prime}-\dots(-1)^{\ell+1}x_{\ell}^{\prime}),0)=(0,0).$
Here we note that the first component of the left-hand side of Equation (5) contains the same number of ++ and - signs. Also, if some element $d=(x,a)$ appears both in both the left-hand and right-hand sides of Equation (4), then we must have $x=x_{k}=x_{b}^{\prime}$ where $k\equiv b\pmod{2}$. Hence $x$ appears in Equation (5) with different signs, so we still have the same number of ++ and - signs after erasing these repeated elements. Thus Equation (5) becomes
$(x_{1}^{\prime\prime}-x_{2}^{\prime\prime}+x_{3}^{\prime\prime}-\dots-x_{t}^{\prime\prime},0)=(0,0).$
Setting $d_{k}^{\prime\prime}=(x_{k}^{\prime\prime},\pm a_{k}^{\prime\prime})$ (according to whether $d_{k}^{\prime\prime}$ comes from the right- or left-hand side of Equation (4)) this implies that
$d_{1}^{\prime\prime}d_{2}^{\prime\prime}\cdots d_{t}^{\prime\prime}=id.$
Since $d_{k}^{\prime-1}=(x_{k}^{\prime},-a_{k}^{\prime})$, the previous relation is in contradiction with the fact that $D$ is dissociated. ∎
4. Ordering the Set $E$ and the Dissociated Sets
Now, the proof proceeds mainly in two steps. Due to the Structure Theorem 3.5, we can assume without loss of generality that
$A=E\cup(\cup_{j=1}^{s}D_{j})$
where the $D_{j}$s are dissociated and satisfy the hypotheses of Theorem 3.5 and we have that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)$. Recall that, given $d\in\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $H$ is abelian and $\varphi$ satisfies property $(\star)$, we let $\pi_{1}(d)$ and $\pi_{2}(d)$ be the projections of $d$ over $\mathbb{Z}_{p}$ and $H$, respectively. We begin by ordering the set $E$ and, for this purpose, set some new notations. We split $E$ into four sets:
$P=\left\{x\in E:\varphi(\pi_{2}(x))=id\mbox{ and }\pi_{1}(x)\in\left(0,\frac{p}{90(|E|+1)}\right)\right\},$
(6) $N=\left\{x\in E:\varphi(\pi_{2}(x))=id\mbox{ and }\pi_{1}(x)\in\left(-\frac{p}{90(|E|+1)},0\right)\right\},$
$Z=\left\{x\in E:\varphi(\pi_{2}(x))=id\mbox{ and }\pi_{1}(x)=0\right\},$
$S=\{x\in E:\varphi(\pi_{2}(x))=-id\}.$
Also, if $S$ is nonempty, we split it into $S=S_{e}\cup S_{o}$ such that $|S_{e}|=\lceil|S|/2\rceil$, $|S_{o}|=\lfloor|S|/2\rfloor$, and
(7) $\forall x\in S_{e},\forall y\in S_{o},\ \pi_{1}(x)\geq\pi_{1}(y).$
Finally, given an ordering $\mathbf{x}=x_{0},x_{1},\dots,x_{m}$, we consider the set of partial products
$\operatorname{IS}(\mathbf{x})=\left\{x_{1}\cdot x_{2}\cdots x_{i}:i\in[0,m]\right\}.$
By the Structure Theorem 3.5, we have that $\delta=(z_{0},a_{0})$ where $\varphi(a_{0})=id$ and $z_{0}\neq 0$.
We consider the case where $S$ is empty in the remark below.
Remark 4.1.
According to Lemma 3.4, up to absorbing, in the proof of Theorem 3.5, one element of $Z$ into the dissociated sets $D_{j}$s, we can impose that $\sum_{z\in Z}z\neq id$.
Now we show that, for any family of subsets $Y_{1},\ldots,Y_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$, we can order $E$ as long as $S\neq\emptyset$ as follows:
Proposition 4.2.
Let $p$ be large enough. Consider $\delta=(z_{0},0)\in\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $z_{0}>0$ and subsets $Y_{1},\ldots,Y_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$. Given a finite subset $E$ such that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$, consider $P,N,Z,S,S_{e}$ and $S_{o}$ defined as above where $S\neq\emptyset$. Then there are orderings $\mathbf{p}$ of $P$, $\mathbf{n}$ of $N$, $\mathbf{z}$ of $Z$, $\mathbf{s}_{e}=(s_{0},s_{2},\dots,s_{2\ell})$ of $S_{e}$ (here $\ell=\lceil|S|/2\rceil-1$) and $\mathbf{s}_{o}=(s_{1},s_{3},\dots,s_{2h-1})$ of $S_{o}$ (here $h=\lfloor|S|/2\rfloor$) such that, setting
(8) $\mathbf{x}=\mathbf{z}\ \mathbf{p}\ s_{0}\ \mathbf{n}\ s_{1}s_{2}\dots,s_{\ell+h},$
we have that $\delta,\mathbf{x}$ is a sequencing and
(9) $|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leq 4\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)+|H|$
for all $1\leq j\leq m$.
Proof.
First, note that since $H$ is strongly sequenceable, there exists a sequencing $\mathbf{z}$ of $Z$. Then for such an ordering we can easily see that any ordering defined in (8) is a sequencing for $E$. Also, since
$\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H,$
we can assume that the first components of $E$ are integers (positive for the elements in $P$ and negative for the elements in $N$).
Now, we will construct the sequences $\mathbf{x}=\mathbf{z}\ \mathbf{p}\ s_{0}\ \mathbf{n}\ s_{1}s_{2}\dots,s_{\ell+h}=\mathbf{z}\ x_{1}x_{2}\dots x_{|P|+|N|+|S|}$ from the larger indices to the smaller indices (with respect to $\mathbf{x}$). We set $x_{i}=(z_{i},\epsilon_{i})$ where $\epsilon_{i}\in\{0,1\}$ according to whether $x_{i}\in P\cup N$ or $x_{i}\in S$, respectively.
Here, we will divide the construction into four cases:
  • 1)
    if $k>|P|+|N|+1$ and k(|P|+|N|+1)k-(|P|+|N|+1) is odd, we need to choose an $x_{k}\in S_{o}$;
  • 2)
    if $k>|P|+|N|+1$ and k(|P|+|N|+1)k-(|P|+|N|+1) is even or if $k=|P|+1$, we need to choose an $x_{k}\in S_{e}$;
  • 3)
    if $|P|+|N|+1\geq k>|P|+1$, we need to choose an $x_{k}\in N$;
  • 4)
    if $|P|+1>k$, we need to choose an $x_{k}\in P$.
After we have chosen $x_{|P|+|N|+|S|},x_{|P|+|N|+|S|-1},\dots,x_{k+1}$, we set:
$S_{o}^{k}:=S_{o}\setminus\{x_{|P|+|N|+|S|},x_{|P|+|N|+|S|-1},\dots,x_{k+1}\},$
$S_{e}^{k}:=S_{e}\setminus\{x_{|P|+|N|+|S|},x_{|P|+|N|+|S|-1},\dots,x_{k+1}\},$
$N^{k}:=N\setminus\{x_{|P|+|N|+|S|},x_{|P|+|N|+|S|-1},\dots,x_{k+1}\},$
$P^{k}:=P\setminus\{x_{|P|+|N|+|S|},x_{|P|+|N|+|S|-1},\dots,x_{k+1}\},$
to be the sets of remaining elements of $S_{o},S_{e},N$ and $P$ and we define the quantity
$\nu_{k}:=\delta+\sum_{x\in P^{k}\cup S_{e}^{k}}x-\sum_{x\in N^{k}\cup S_{o}^{k}}x.$
Note that, for any ordering $\mathbf{x}=\mathbf{z}\ x_{1}x_{2}\dots x_{|P|+|N|+|S|}$, we have that $\nu_{k}=\delta\cdot\prod_{z\in Z}z\cdot\prod_{i=1}^{k}x_{i}$.
Suppose now we are in the first case. This means that we have already fixed
$x_{|P|+|N|+|S|},x_{|P|+|N|+|S|-1},\dots,x_{k+1}$
alternatively in $S_{e}$ and in $S_{o}$ (according to the parity of k(|P|+|N|+1)k-(|P|+|N|+1)) and we need to choose $x_{k}\in S_{o}$. Following the idea of Proposition 4.1 of [4], we choose $x_{k}$ as follows:
  1. (1)
    If there is some $x^{*}\in S_{o}^{k}$ such that $\nu_{k}+x^{*}\notin\cup_{j}Y_{j}$, then choose $x_{k}$ to be this $x^{*}$ and say that the current step is a skip-step for $S_{o}$.
  2. (2)
    Now, consider the case where $\nu_{k}+S_{o}^{k}\subseteq\cup_{j}Y_{j}$. Let $i$ be minimal such that $\nu_{k}+S_{o}^{k}$ intersects $Y_{i}$, and say that the current step is an $i$-step for $S_{o}$. If $\nu_{k}+S_{o}^{k}\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\min\{\pi_{1}(x):x\in S_{o}^{k}\}.$
    If $\nu_{k}+S_{o}^{k}\not\subseteq Y_{i}$, then we set
    $x_{k}:\pi_{1}(x_{k})=\min\{\pi_{1}(x):x\in S_{o}^{k}\mbox{ and }\nu_{k}+x\not\in Y_{i}\}.$
In the second, third, and fourth cases, we choose $x_{k}\in S^{k}_{e},N^{k},P^{k}$, respectively, by the same method as in case 1 where $x_{k}$ is chosen as a minimum in the case of $N^{k}$ and as a maximum otherwise.
Now we prove that the ordering $\mathbf{x}$ satisfies
$|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leq 4\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)+|H|$
for all $1\leq j\leq m$.
We split the set $\delta\cdot\operatorname{IS}(\mathbf{x})$ into the following subsets:
$\delta\cdot\operatorname{IS}^{o}(\mathbf{x})=\{\nu_{\ell}:\ell\in[|N|+|P|+2,|N|+|P|+|S|]\mbox{ and }\ell-(|N|+|P|+|S|)\mbox{ is odd}\},$
$\delta\cdot\operatorname{IS}^{e}(\mathbf{x})=\{\nu_{\ell}:\ell\in[|N|+|P|+2,|N|+|P|+|S|]\mbox{ and }\ell-(|N|+|P|+|S|)\mbox{ is even}\}$
$\cup\{(\nu_{\ell},1):\ell=|P|+1\},$
$\delta\cdot\operatorname{IS}^{p}(\mathbf{x})=\{\nu_{\ell}:\ell\in[0,|P|]\},$
$\delta\cdot\operatorname{IS}^{n}(\mathbf{x})=\{\nu_{\ell}:\ell\in[|P|+2,|N|+|P|+1]\}.$
Equation (9) follows from $|\delta\cdot\operatorname{IS}(\mathbf{z})|\leq|H|$ and by showing that $|\delta\cdot\operatorname{IS}^{p}(\mathbf{x})\cap Y_{j}|$, $|\delta\cdot\operatorname{IS}^{n}(\mathbf{x})\cap Y_{j}|$, $|\delta\cdot\operatorname{IS}^{e}(\mathbf{x})\cap Y_{j}|$, $|\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}|$ are upper bounded by $\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)$. We will prove the statement of Equation (9) (without the factor of four) only for $|\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}|$ since the argument for $|\delta\cdot\operatorname{IS}^{e}(\mathbf{x})\cap Y_{j}|$ is essentially identical and that of $|\delta\cdot\operatorname{IS}^{p}(\mathbf{x})\cap Y_{j}|$ and $|\delta\cdot\operatorname{IS}^{n}(\mathbf{x})\cap Y_{j}|$ is identical to that of [4].
Given some value of $\ell$, $\nu_{\ell}$ lie in $Y_{j}$ only when the choice of $x_{\ell+1}$ is a $j$-step or an $i$-step (for $S_{e}$) for some $i<j$. We will estimate these two contributions separately. Note that skip-steps and $i$-steps for $i>j$ never contribute.
We first consider the contribution of $j$-steps. Suppose that the choice of $x_{\ell+1}=(z_{\ell+1},a_{\ell+1})$ is a $j$-step for $S_{e}$ and $\nu_{\ell}=\nu_{\ell+1}-x_{\ell+1}\in Y_{j}$. Then we must have
$\nu_{\ell+1}-S_{e}^{\ell+1}\subseteq Y_{j}.$
Since $z_{\ell+1}$ is maximal in $\pi_{1}(S_{e}^{\ell+1})$, given $z^{*}\in\pi_{1}(S_{e}^{\ell+1})$, we have that
$\pi_{1}(\nu_{\ell+1})-z^{*}\geq\pi_{1}(\nu_{\ell+1})-z_{\ell+1}=\pi_{1}(\nu_{\ell}).$
Moreover,
$\pi_{1}(\nu_{\ell+2})=\pi_{1}(\nu_{\ell+1})-\pi_{1}(x_{\ell+2})=(\pi_{1}(\nu_{\ell+1})-z^{*})+(z^{*}-\pi_{1}(x_{\ell+2}))$
and hence due to Equation (7) we have
$\pi_{1}(\nu_{\ell+2})\geq\pi_{1}(\nu_{\ell+1})-z^{*}\geq\pi_{1}(\nu_{\ell}).$
Due to the monotonicity of $\pi_{1}(\nu_{|P|+|N|+1+2t+1})$, setting $\ell=|P|+|N|+1+2t+1$, it follows that the (t1)/|H|(t-1)/|H| elements of $\nu_{\ell+1}-S_{e}^{\ell+1}\setminus\{\nu_{\ell}\}$ can never appear in $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})$. In particular, from such $j$-steps with $t\geq L+1$ we obtain at most $|H||Y_{j}|/L$ elements of $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}$. From $j$-steps with $t\leq L$ we trivially obtain at most $L$ elements of $\delta\cdot\operatorname{IS}^{o}(\mathbf{x})\cap Y_{j}$.
We now consider the contribution of $i$-steps for $S_{e}$ with $i<j$. We will bound this contribution by the total number of $i$-steps for $S_{e}$ with $i<j$. We claim that the number of $i$-steps is at most $2|H||Y_{i}|$ for each $i$. For each $i$-step $\ell+1$, let $y(\ell)$ denote the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1}\cap Y_{i})$. It suffices to show that each $y\in\pi_{1}(Y_{i})$ appears as $y(\ell)$ for at most $2|H|$ different $i$-steps (for $S_{e}$).
If $y(\ell)$ is not the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1})$, then it is distinct from $y(\ell^{\prime})$ for all odd $\ell^{\prime}<\ell$ since
$y(\ell)>\pi_{1}(\nu_{\ell+1})-z_{\ell+1}=\pi_{1}(\nu_{\ell})$
and, given $z^{\prime}\in\pi_{1}(S_{e}^{\ell-1})$, by Equation (7) we have
$\pi_{1}(\nu_{\ell})\geq\pi_{1}(\nu_{\ell})+(z_{\ell}-z^{\prime})=\pi_{1}(\nu_{\ell-1})-z^{\prime}\geq y(\ell-2).$
Finally, set $\ell=|P|+|N|+1+2t+1$ and note that the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1})$ is nondecreasing as $t$ increases. It follows that $y(\ell)$ is the smallest element of $\pi_{1}(\nu_{\ell+1})-\pi_{1}(S_{e}^{\ell+1})$ for at most $|H|$ values of $\ell$. Summing up, we obtain that
$|\delta\cdot\operatorname{IS}(\mathbf{x})\cap Y_{j}|\leq 4\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)+|H|$
for all $1\leq j\leq m$. Since such an inequality holds also for $\delta\cdot\operatorname{IS}^{e}(\mathbf{x})$, $\delta\cdot\operatorname{IS}^{p}(\mathbf{x})$, and $\delta\cdot\operatorname{IS}^{n}(\mathbf{x})$, we get Equation (9). ∎
With essentially the same proof of Proposition 4.2 and thanks to Claims 4.2 and 4.3 of [4] we obtain the following proposition which shows that, for any family of subsets $Y_{1},\ldots,Y_{m},Y^{\prime}_{1},\ldots,Y^{\prime}_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$, we can order $E$ in the case of $S=\emptyset$ as follows:
Proposition 4.3.
Let $p$ be large enough. Consider $\delta=(z_{0},0)\in\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $z_{0}>0$ and subsets $Y_{1},\ldots,Y_{m},Y^{\prime}_{1},\ldots,Y^{\prime}_{m}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$. Take a finite subset $E$ such that $\pi_{1}(E)\subseteq\left(-\frac{p}{90(|E|+1)},\frac{p}{90(|E|+1)}\right)\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ and consider $P,N,Z,S$ to be defined as above where $S=\emptyset$. Then there are orderings $\mathbf{p}$ of $P$, $\mathbf{n}$ of $N$, and $\mathbf{z}$ of $Z$ such that $\mathbf{z},\overline{\mathbf{p}},\delta,\mathbf{n}$ is a sequencing and
(10) $|\delta\cdot\operatorname{IS}(\mathbf{p})\cap Y_{j}|\leq\inf_{L\in\mathbb{N}}\left(\frac{|H||Y_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y_{i}|\right)$
(11) $|\delta\cdot\operatorname{IS}(\mathbf{n})\cap Y^{\prime}_{j}|\leq\inf_{L\in\mathbb{N}}\left(\frac{|H||Y^{\prime}_{j}|}{L}+L+2|H|\sum_{i=1}^{j-1}|Y^{\prime}_{i}|\right)$
for all $1\leq j\leq m$.
Now the proof proceeds by ordering the dissociated sets via a probabilistic procedure inspired by that of [4]. In particular, Lemma 5.1 of [4] allows us to split a dissociated set $D$, where the $H$ component of each element of $D$ is in $H_{0}$, into uniform random partitions $D^{(1)},D^{(2)},D^{(3)}$, and $D^{(4)}$. Consequently, we need only consider the case of dissociated sets $D$ with $H$ components in $H_{1}$. Note that we use the notation $\Omega$ as defined in [4].
Lemma 4.4.
Let $D\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ be a dissociated set with $H$ components in $H_{1}$ and where $|D|$ is divisible by $8$. We split $D$ into $D^{o}=\{d_{1}^{o},d_{2}^{o}\dots,d_{h}^{o}\}$ and $D^{e}=\{d_{1}^{e},d_{2}^{e}\dots,d_{h}^{e}\}$ where $h=|D^{o}|=|D^{e}|=\frac{1}{2}|D|$. Then we let $D^{o}=D^{o,(1)}\cup D^{o,(2)}\cup D^{o,(3)}\cup D^{o,(4)}$ and $D^{e}=D^{e,(1)}\cup D^{e,(2)}\cup D^{e,(3)}\cup D^{e,(4)}$ be uniformly random partitions of $D^{o}$ and $D^{e}$, where each partition has the same cardinality. Then for every nonempty proper interval $I\subseteq\{1,2,3,4\}$ and each $x\in\mathbb{Z}_{p}\rtimes_{\varphi}H$, we have
$\mathbb{P}\left(\prod_{i\in I}\prod_{\begin{subarray}{c}d^{o}\in D^{o,(i)}\\ d^{e}\in D^{e,(i)}\end{subarray}}d^{o}\cdot d^{e}=x\right)\leq e^{-\Omega(|D|)}$
Proof.
Choose uniformly random partition $D^{e}=D^{e,(1)}\cup D^{e,(2)}\cup D^{e,(3)}\cup D^{e,(4)}$ where $|D^{e,(1)}|=|D^{e,(2)}|=|D^{e,(3)}|=|D^{e,(4)}|=\frac{|D|}{8}$ is an integer by assumption. Take a similar partition for $D^{o}$. For each $i\in\{1,2,3,4\}$, the $\binom{|D|/2}{|D|/8}^{2}$ possible values of
$\prod_{\begin{subarray}{c}d^{o}\in D^{o,(i)}\\ d^{e}\in D^{e,(i)}\end{subarray}}d^{o}\cdot d^{e}$
are achieved with equal probability, as are the $\binom{|D|/2}{|D|/4}^{2}$ possible values of
$\prod_{\begin{subarray}{c}d^{o}\in D^{o,(i)}\cup D^{o,(j)}\\ d^{e}\in D^{e,(i)}\cup D^{e,(j)}\end{subarray}}d^{o}\cdot d^{e}$
and the $\binom{|D|/2}{3|D|/8}^{2}$ possible values of
$\prod_{\begin{subarray}{c}d^{o}\in D^{o,(i)}\cup D^{o,(j)}\cup D^{o,(k)}\\ d^{e}\in D^{e,(i)}\cup D^{e,(j)}\cup D^{e,(k)}\end{subarray}}d^{o}\cdot d^{e}.$
Since $\binom{|D|/2}{|D|/8},\binom{|D|/2}{|D|/4}$, and $\binom{|D|/2}{3|D|/8}$ are all $e^{\Omega(|D|)}$, their squares must also be $e^{\Omega(|D|)}$. The statement of the lemma follows from this. ∎
As in [4], we reorder the dissociated sets as follows. For each dissociated set $D_{j}$ with $j\in L_{1}$, we randomly partition $D_{j}=\cup_{i=1}^{4}D_{j}^{(i)}$ into four sets of equal size, where each $D_{j}^{(i)}=D_{j}^{o,(i)}\cup D_{j}^{e,(i)}$ as described in Lemma 4.4. For each dissociated set $D_{j}$ with $j\in L_{0}$, we operate analogously to Lemma 5.1 of [4] in that we partition uniformly at random $D_{j}=\cup_{i=1}^{4}D_{j}^{(i)}$ into four sets of equal size. All these partitions are performed independently. Then we order these new dissociated sets as follows:
(12) $D_{1}^{(1)},D_{1}^{(2)},D_{2}^{(1)},D_{2}^{(2)},\ldots,D_{s}^{(1)},D_{s}^{(2)},D_{1}^{(3)},D_{1}^{(4)},D_{2}^{(3)},D_{2}^{(4)},\ldots,D_{s}^{(3)},D_{s}^{(4)}\,.$
For notational convenience, we rename this new sequence as $T_{1},T_{2},\ldots,T_{u}$ where $u=4s$, and we use the same notation $T_{j}^{o}$ and $T_{j}^{e}$ as used for the $D_{j}$s. We define $\tau_{j}:=\prod_{d^{o}\in T_{j}^{o},d^{e}\in T_{j}^{e}}d^{o}\cdot d^{e}$ if $T_{j}$ is a dissociated set associated to a $D_{k}$ with $k\in L_{1}$, otherwise we define it as $\tau_{j}:=\prod_{d\in T_{j}}d$. We use $\mathbf{t}_{i}$ to denote an ordering of $T_{i}$ that is chosen according to a suitable distribution. Let us also define
(13) $\prod_{\leq m}(T_{j}):=\left\{\prod_{d^{o}\in S^{o},d^{e}\in S^{e}}d^{o}\cdot d^{e}:S^{o}\subseteq T_{j}^{o},S^{e}\subseteq T_{j}^{e},|S^{o}|=\lceil k/2\rceil,|S^{e}|=\lfloor k/2\rfloor,k\leq m\right\}$
if $T_{j}$ is a dissociated set associated to a $D_{k}$ with $k\in L_{1}$. Otherwise, we define it as
(14) $\prod_{\leq m}(T_{j}):=\left\{\prod_{d\in S}d:S\subseteq T_{j},|S|\leq m\right\}\,.$
Given a subset $S$ of a finite group, we define $(S)^{-1}$ to be the set $\{d^{-1}:d\in S\}$.
Recalling the definition given in [4] of Type I and Type II intervals, we say that a proper nonempty interval $I\subset[1,|A|]$ is of Type II if it contains between $K$ and $|T_{j}|-K$ elements of some $T_{j}$, where $K=c_{2}R^{1/3}$ for a sufficiently small positive real $c_{2}$; otherwise we say that the interval is of Type I.
In the following lemma, we adapt Lemma 5.3 of [4] to the semidirect product case.
Lemma 4.5.
Let $p$ be large enough and let $c>0$ and $1\leq s\leq e^{c(\log p)^{1/4}}$. Take a sequence of dissociated sets $D_{1},D_{2},\ldots,D_{s}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ each of size $\asymp R$ and a multiple of eight, where $H$ is abelian and $\varphi$ satisfies property $(\star)$. Let $\mathbf{x}_{1},\mathbf{x}_{2}$ be sequences over $\mathbb{Z}_{p}\rtimes_{\varphi}H$ of length at most $e^{c(\log p)^{1/4}}$, and assume that $\mathbf{x}_{1},\delta,\mathbf{x}_{2}$ is a sequencing. If the sequence $T_{1},\ldots,T_{u}$ of dissociated sets is chosen randomly as described previously, then we have with positive probability that
  1. $(i)$
    for each proper nonempty interval $I=[i,j]\subseteq[1,u]$, we have that
    $id\not\in\left(\prod_{\leq K}(T_{i-1})\cup\left(\prod_{\leq K}(T_{i})\right)^{-1}\right)\cdot\tau_{i}\cdots\tau_{j}\cdot\left(\left(\prod_{\leq K}(T_{j})\right)^{-1}\cup\prod_{\leq K}(T_{j+1})\right)$
    with the convention that $T_{0}=T_{u+1}=\emptyset$
  2. $(ii)$
    for each $2\leq j\leq u$, we have that
    $id\not\in\operatorname{IS}(\mathbf{x}_{1})\cdot\tau_{1}\cdots\tau_{j}\cdot\left(\left(\prod_{\leq K}(T_{j})\right)^{-1}\cup\prod_{\leq K}(T_{j+1})\right)$
    $id\not\in\operatorname{IS}(\mathbf{x}_{2})\cdot\tau_{u}\cdots\tau_{j}\cdot\left(\left(\prod_{\leq K}(T_{j})\right)^{-1}\cup\prod_{\leq K}(T_{j-1})\right)$
  3. $(iii)$
    the ordering $\mathbf{x}_{1},\tau_{1},\ldots,\tau_{u},\mathbf{x}_{2}$ is a sequencing
Proof.
Thanks to Lemma 4.4 and Lemma 5.1 of [4], we have that $\mathbb{P}(\prod_{i\in I}\tau_{i}=x)\leq e^{-\Omega(R)}$ for any proper nonempty subinterval $I\subset[1,u]$ and any element $x\in\mathbb{Z}_{p}\rtimes_{\varphi}H$. Observe that $(iii)$ holds whenever $(i)$ and $(ii)$ hold since we have assumed that $\mathbf{x}_{1},\delta,\mathbf{x}_{2}$ is a sequencing. Then since $|T_{j}|\asymp R$ and for $K=o(R)$ we have $|\prod_{\leq K}(T_{j})|=e^{o(R)}$, it is clear that $(i)$ fails for a subinterval $I$ with probability at most $e^{-\Omega(R)}$. Hence, by the union bound over all the proper nonempty subintervals of $[1,u]$, it can be seen that the condition $(i)$ fails with probability $o(1)$ as $p\to\infty$. The proof of $(ii)$ can be easily derived by noticing that $|\operatorname{IS}(\mathbf{x}_{1})|,|\operatorname{IS}(\mathbf{x}_{2})|\leq e^{c(\log p)^{1/4}}$. ∎
We note that Lemma 4.5 handles all Type I intervals except the ones starting in the first $K$ elements of $\mathbf{t}_{2j}$ and ending in the last $K$ elements of $\mathbf{t}_{2j+1}$ and the ones starting in $\mathbf{x}$ and ending in the first $K$ elements of $\mathbf{t}_{1}$ or in the last $K$ elements of $\mathbf{t}_{u}$.
From the definitions in Equations (13) and (14) we can derive in a natural way the definitions of $\prod_{=m}(T_{j})$. Then, we can state the following lemma, which is an adaptation of Lemma 5.4 of [4].
Lemma 4.6.
Let $T_{1}$ and $T_{u}$ be the two subsets chosen randomly as described previously. Then with positive probability we have for $h=1,u$, that
(15) $\left|\prod_{=j}(T_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\leq 4K\frac{\binom{|T_{j}^{o}|}{\lceil j/2\rceil}\binom{|T_{j}^{e}|}{\lfloor j/2\rfloor}}{\binom{|D_{j}^{o}|}{\lceil j/2\rceil}\binom{|D_{j}^{e}|}{\lfloor j/2\rfloor}}\left|\prod_{=j}(D_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\,$
if $T_{h}$ is a dissociated set whose $H$ components are in $H_{1}$. Otherwise
(16) $\left|\prod_{=j}(T_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\leq 4K\frac{\binom{|T_{j}|}{j}}{\binom{|D_{j}|}{j}}\left|\prod_{=j}(D_{h})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|\,.$
Proof.
Here $T_{1}$ is chosen uniformly at random from all the subsets of $D_{1}$ with size equal to $|D_{1}|/4$ if $D_{1}$ is a dissociated set whose $H$ components are in $H_{1}$. Otherwise, $T_{1}$ is the union of two sets $T_{1}^{o}$ and $T_{1}^{e}$ where each of these sets is chosen uniformly at random from all the subsets of $D_{1}^{o}$ and $D_{1}^{e}$, respectively, with size equal to $|D_{1}|/8$. One can see that the expected value of the left-hand side of inequalities (15) and (16) is exactly the right-hand side of those equations without the factor $4K$. Therefore, by Markov’s inequality, we see that bounds (15) and (16) fail with probability at most $1/4K$. The same argument applies also for $T_{u}$. Finally, we conclude the proof of this lemma by taking the union bound over all $j\in[1,K]$ and $h=1,u$. ∎
Now, recalling that $K=c_{2}R^{1/3}$, we set
(17) $Y_{j}:=\left(\prod_{=j}(D_{s})\right)^{-1}\cup\left(\delta^{-1}\prod_{=j}(D_{1})\right)$
(18) $Y^{\prime}_{j}:=\left(\prod_{=j}(D_{1})\right)^{-1}\cup\left(\delta^{-1}\prod_{=j}(D_{s})\right)$
for each $1\leq j\leq K$ and we apply Proposition 4.2. This gives us an ordering of $\mathbf{x}$ such that $\delta,\mathbf{x}$ is a sequencing, and inequality (9) holds. Then, using Lemmas 4.5 and 4.6, we obtain dissociated sets $T_{1},\ldots,T_{u}$ from $D_{1},\ldots,D_{s}$ such that $A=P\cup N\cup Z\cup S\cup\left(\cup_{i=1}^{u}T_{i}\right)$ satisfy the conditions imposed in these lemmas. Fixing such a choice for $T_{1},\ldots,T_{u}$, we will construct a sequencing of $A$ of the form $\mathbf{x}_{1},\mathbf{t}_{1},\ldots,\mathbf{t}_{u},\mathbf{x}_{2}$, where the $\mathbf{t}_{i}$s are random orderings chosen with a suitable distribution described later. Now, we take care of the remaining Type I intervals left open by Lemmas 4.5 and 4.6. Following the procedure of [4], we pick the orders of each $\mathbf{t}_{i}$ uniformly at random on a restricted subset of the possible orderings.
We say that an ordering $t_{1},t_{2},\ldots,t_{|T_{1}|}$ of $T_{1}$ is acceptable if
$t_{1}\cdot t_{2}\cdots t_{k}\not\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\,$
for every $1\leq k\leq K$, and we say that an ordering $t_{1},t_{2},\ldots,t_{|T_{u}|}$ of $T_{u}$ is acceptable if
$t_{1}\cdot t_{2}\cdots t_{k}\not\in-\operatorname{IS}(\mathbf{x}_{2})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{1})\right)\,$
for every $1\leq k\leq K$. We can now state the following lemma, which is an adaptation of Lemma 6.1 of [4].
Lemma 4.7.
If $T_{1}$ is a dissociated set whose $H$ components are in $H_{1}$, pick an ordering of $T_{1}$ uniformly at random from the orderings of $T_{1}$; otherwise pick the elements in odd positions of such an ordering uniformly at random from the orderings of $T_{1}^{o}$ and the elements in even positions uniformly at random from the orderings of $T_{1}^{e}$. Pick a random ordering of $T_{u}$ by the same process. Then these two orderings are acceptable with positive probability.
Proof.
We prove the statement only for $T_{1}$ since the proof is analogous for $T_{u}$. Let $\mathbf{t}_{1}=t_{1},\ldots,t_{|T_{1}|}$ be a random ordering of $T_{1}$. If $T_{1}$ is a dissociated set whose $H$ components are in $H_{1}$ then $\mathbf{t}_{1}$ is chosen uniformly at random from the orderings of $T_{1}$; otherwise $t_{1},t_{3},\ldots$ are chosen uniformly at random from the orderings of $T_{1}^{o}$ and $t_{2},t_{4},\ldots$ from the orderings of $T_{1}^{e}$. We want to prove that
$\mathbb{P}\left(t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\leq\frac{1}{100K}\,$
for each $1\leq k\leq K$. For a fixed value of $k$, since $T_{1}$ is a dissociated set, Lemma 3.6 tells us that the products $t_{1}\cdots t_{k}$ are uniformly distributed over the set $\prod_{=k}(T_{1})$. Given the definition of the sets $Y_{j}$s in (17), we see that $|Y_{j}|=2\left|\prod_{=j}(D_{1})\right|$ since $|D_{1}|=|D_{s}|$ and $D_{1}\cup D_{s}\cup\{\delta\}$ is dissociated by condition $(v)$ of the Structure Theorem. Then, by Proposition 4.2, taking $L=\lfloor|H|^{1/2}|Y_{k}|^{1/2}\rfloor$ and $K=c_{2}R^{1/3}$, we have that
$\left|\prod_{=k}(D_{1})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|$ $=O\left(|H|^{1/2}|Y_{k}|^{1/2}+|H|\sum_{i=1}^{k-1}|Y_{i}|+|H|\right)$
$=O\left(\left|\prod_{=k}(D_{1})\right|\cdot\frac{K}{|D_{1}|}\cdot|H|\right)$
(19) $=O\left(\left|\prod_{=k}(D_{1})\right|\cdot|H|\cdot c_{2}^{3}K^{-2}\right)\,.$
Since the right-hand side of Equation (16) is an upper bound for the right-hand side of Equation (15), by Lemma 4.6 and Equation (4) we obtain
$\left|\prod_{=k}(T_{1})\cap\left(-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\right|=O\left(4K\left|\prod_{=k}(T_{1})\right|\cdot|H|\cdot c_{2}^{3}K^{-2}\right)=O\left(\left|\prod_{=k}(T_{1})\right|\cdot|H|\cdot c_{2}^{3}K^{-1}\right)\,.$
Then, it follows that
$\mathbb{P}\left(t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)=O\left(|H|\cdot c_{2}^{3}K^{-1}\right)\,.$
Therefore, for a sufficiently small real $c_{2}$ that depends on $|H|$, we have that $\mathbb{P}\left(t_{1}\cdots t_{k}\in\delta\cdot\operatorname{IS}(\mathbf{x})\right)\leq 1/(100K)$. The theorem follows by the union bound over $1\leq k\leq K$ since
$\mathbb{P}\left(\cup_{k=1}^{K}t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\leq\sum_{k=1}^{K}\mathbb{P}\left(t_{1}\cdots t_{k}\in-\operatorname{IS}(\mathbf{x}_{1})\cup\left(\delta\cdot\operatorname{IS}(\mathbf{x}_{2})\right)\right)\leq\frac{1}{100}\,.$
Following [4], we choose $\mathbf{t}_{1},\mathbf{t}_{u}$ independently such that $\mathbf{t}_{1},\mathbf{t}_{u}$ are acceptable orderings (here $\overline{\mathbf{t}}_{u}=t_{|T_{1}|},\ldots,t_{1}$ denotes the reverse of $\mathbf{t}_{u}$). This choice takes care of Type I intervals with one endpoint in $\mathbf{x}$ and the other one in the first $K$ elements of $\mathbf{t}_{1}$ or in the last $K$ elements of $\mathbf{t}_{u}$. Now, we choose the pair of orderings $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ for $j\in[1,u/2-1]$ as follows: given a partial ordering $t_{1},\ldots,t_{k}$ of $T_{2j}$ where if $T_{2j}$ is a dissociated set whose $H$ components are in $H_{1}$ then $t_{1},t_{3},\ldots$ is a partial ordering of $T_{2j}^{o}$ and $t_{2},t_{4},\ldots$ is a partial ordering of $T_{2j}^{e}$, and given a partial ordering $t^{\prime}_{1},\ldots,t^{\prime}_{\ell}$ of $T_{2j+1}$ by the same process, we say that the pair of partial orderings is permissible if $t_{1}\cdots t_{i}\cdot t^{\prime}_{1}\cdots t^{\prime}_{s}\neq id$ for all $(i,s)$. We let $t_{1},t_{2},\ldots,t_{K}$ and $t^{\prime}_{1},t^{\prime}_{2},\ldots,t^{\prime}_{K}$ be a uniformly random permissible pair of orderings, then we choose $\mathbf{t}_{2j}$ to be a uniformly random ordering of $T_{2j}$ if $T_{2j}$ is a dissociated set whose $H$ components are in $H_{0}$. Otherwise, the elements in odd positions are picked uniformly at random from $T_{2j}^{o}$ and the elements in even positions from the orderings of $T_{2j}^{e}$ conditional on $\overline{\mathbf{t}}_{2j}$ starting with $t_{1},\ldots,t_{K}$. Moreover, $\mathbf{t}_{2j+1}$ is picked with the same process conditional on $\mathbf{t}_{2j+1}$ starting with $t^{\prime}_{1},\ldots,t^{\prime}_{K}$. These choices are made independently for each $j$ and from the choices of $\mathbf{t}_{1},\mathbf{t}_{u}$.
As done in [4], we can see that also in our case, for any $1\leq j\leq u/2-1$ and any orderings of $\mathbf{t}_{2j}$ and $\mathbf{t}_{2j+1}$, we have that
$\mathbb{P}\left(\mathbf{t}_{2j+1}=t^{\prime}_{1},\ldots,t^{\prime}_{|T_{2j+1}|}\mid\overline{\mathbf{t}}_{2j}=t_{1},\ldots,t_{|T_{2j}|}\right)=O\left(\frac{1}{|T_{2j+1}|!}\right)$
(20) $\mathbb{P}\left(\mathbf{t}_{2j}=t_{1},\ldots,t_{|T_{2j}|}\mid\mathbf{t}_{2j+1}=t^{\prime}_{1},\ldots,t^{\prime}_{|T_{2j+1}|}\right)=O\left(\frac{1}{|T_{2j}|!}\right)\,.$
We note that this choice on the pairs $\mathbf{t}_{2j},\mathbf{t}_{2j+1}$ handles Type I intervals starting in the first $K$ elements of $\mathbf{t}_{2j}$ and ending in the last $K$ elements of $\mathbf{t}_{2j+1}$.
Now, if we choose randomly the orderings $\mathbf{t}_{1},\mathbf{t}_{2},\ldots,\mathbf{t}_{u}$ as described above, proceeding as in Lemma 6.3 of [4], we have with high probability that the ordering
$\mathbf{x}_{1},\mathbf{t}_{1},\mathbf{t}_{2},\ldots,\mathbf{t}_{u},\mathbf{x}_{2}\,$,
is a sequencing whenever $|\mathbf{x}_{1}|,|\mathbf{x}_{2}|\leq e^{c(\log p)^{1/4}}$ and $p$ is sufficiently large. More precisely, we can state the following theorem:
Theorem 4.8.
Let $H$ be a finite abelian and strongly sequenceable group. Let $c>0$ and let us consider $\mathbb{Z}_{p}\rtimes_{\varphi}H$ where $\varphi$ satisfies property $(\star)$ and $p$ is large enough depending on $H$. Let $D_{1},D_{2},\ldots,D_{s}\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H$ be dissociated sets, each of size $\asymp R$ and a multiple of eight where $1\leq s\leq e^{c(\log p)^{1/4}}$.
Defining $\delta$ as in Theorem 3.5(iv), we assume that $D_{1},D_{s}$, and $\delta$ satisfy condition $(v)$ of that theorem. Now we consider sequences $\mathbf{x}_{1},\mathbf{x}_{2}$ over $\mathbb{Z}_{p}\rtimes_{\varphi}H$ of length at most $e^{c(\log p)^{1/4}}$ such that $\mathbf{x}_{1},\delta,\mathbf{x}_{2}$ is a sequencing satisfying Proposition 4.2 when $\mathbf{x}_{1}=\emptyset$ or Proposition 4.3 when $\mathbf{x}_{1}\neq\emptyset$ with respect to the sets $Y_{j}$s defined in Equation (17) for $\mathbf{x}_{1}=\emptyset$ and with respect to the sets $Y_{j}$s and $Y^{\prime}_{j}$s defined in Equations (17) and (18) for $\mathbf{x}_{1}\neq\emptyset$.
Then it is possible to choose the sequence $T_{1},\ldots,T_{u}$ of dissociated sets and the orderings $\mathbf{t}_{1},\dots,\mathbf{t}_{u}$ so that each $\mathbf{t}_{i}$ is a sequencing and so is
$\mathbf{x}_{1},\mathbf{t}_{1},\ldots,\mathbf{t}_{u},\mathbf{x}_{2}.$
Now, we obtain the main result of our work.
Theorem 1.3.
Let $H$ be a finite abelian and strongly sequenceable group. There exists a $c>0$ such that every subset $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$, where $\varphi:H\to Aut(\mathbb{Z}_{p})$ satisfies $\varphi(h)\in\{id,-id\}$ for each $h\in H$, of size
$|A|\leq e^{c(\log p)^{1/4}},$
has a sequencing.
Proof.
We assume, up to automorphism of $\mathbb{Z}_{p}\rtimes_{\varphi}H$, that
$A=E\cup(\cup_{j=1}^{s}D_{j}),$
where the sets $D_{j}$s and $E$ are defined according to Theorem 3.5. If $A=E$ (i.e. $s=0$) then Theorem 1.3 follows immediately from Proposition 4.2. Otherwise, we have $s>0$ and, up to changing all the signs of the $\mathbb{Z}_{p}$ components, we can assume that $z_{0}>0$ and $\varphi(a_{0})=id$ for the $\delta=(z_{0},a_{0})$ defined in Theorem 3.5(iv). As done in Equation (6), we can split the set $E$ into four sets $P,N,Z$ and $S$. Now, if $S\neq\emptyset$ then we apply Proposition 4.2 to obtain an ordering $\mathbf{x}$ of $E$ such that $\delta,\mathbf{x}$ is a sequencing satisfying Condition (9) with respect to the sets $Y_{j}$s defined in Equation (17), while for the case $S=\emptyset$, we apply Proposition 4.3 to obtain orderings $\mathbf{z}$ of $Z$, $\mathbf{p}$ of $P$, and $\mathbf{n}$ of $N$ such that $\mathbf{z},\overline{\mathbf{p}},\delta,\mathbf{n}$ is a sequencing satisfying Conditions (10) and (11) with respect to the sets $Y_{j}$s and $Y^{\prime}_{j}$s defined in Equations (17) and (18).
Then, let $d_{1}>0$ and $\bar{p}$ such that Theorem 4.8 holds for $c=d_{1}$ and every $p\geq\bar{p}$. Hence for all $p\geq\bar{p}$, we have that all the subsets $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$ of size $|A|\leq e^{d_{1}(\log p)^{1/4}}$ are sequenceable. For $p<\bar{p}$, we set $d_{2}=\log 2/(\log\bar{p})^{1/4}$ to obtain the upper bound $|A|\leq e^{d_{2}(\log p)^{1/4}}<2$. Clearly, all the subsets $A\subseteq\mathbb{Z}_{p}\rtimes_{\varphi}H\setminus\{id\}$ of size at most $1$ are sequenceable. Hence, the statement of the theorem follows by taking $c=\min\{d_{1},d_{2}\}$. ∎
5. Acknowledgements
The first and second authors were partially supported by INdAM–GNSAGA. The authors would also like to thank Noah Kravitz and Benjamin Bedert for useful discussions on this topic.